1. Find the determinant of the follow ing matrices 1 8 [2 1 B=3 8 10 det BT
Answer #1

Al 468) – 1(9) = 23 2) 2 1 - 2 13 % 181=2 (-8-0) -1 (3-4) -2 (0-8) - 16 47+16 (16l = 7 3 BT 5 2 1 1 1-2 3 17 8 0 uri IBTI = 2
Answer #2

source: cofactor expansion of a determinant
answered by: Tom Andy